During a science experiment, Claire created an 18% alcohol solution by mixing 4 fluid ounces of a 28% alcohol solution with a certain amount of a 14% alcohol solution. How many fluid ounces of the 18% alcohol solution did she create?
[A] 12 fl oz [B] 14 fl oz [C] 10 fl oz [D] 18 fl oz

Answer by Boreal(15235)

4 oz*0.28=1.12 oz pure alcohol
x oz* 0.14=0.14 x pure
4+x oz *0.18=0.72+0.18x pure
so 1.12+0.14x=0.72+0.18x
0.40=0.04x
4x=40, multiplying by 100
x=10 oz 18%
C